In the agricultural market for wheat, the demand function is represented by Qd ■ = 80 - 2p and the supply function is Q0, = 40 + 3p. Show what effect the policy of setting a minimum price of S / 10 can have. Determine the equilibrium price and quantity. This measure is cheaper for the government and more profitable for the consumer.
Answer #1

Sol :

Equilibrium is the point where quantity demanded and quantity supplied are equal to each other.

Qd = Qs

80 - 2P = 40 + 3P

40 = 5P

8 = Equilibrium Price

Qd = Qs = 64

the policy of setting minimum price of 10 will have the following effects :

Qd = 80 - 2P

= 80 - 20 = 60

Qs = 40 + 3(10) = 70

Surplus = Qs - Qd

= 70 - 60 = 10

This policy will make the consumer more worse than before as they have to pay higher prices for the wheat . Whereas government have to buy the surplus from the producer at higher prices.